Hi
30 books and magazines were on a shelf. Each book costs $5 while each magazine costs $3. The total cost of the books was $6 more than the magazines.
How many books were on the shelf.
Thanks

Answer by ikleyn(52781)

.
30 books and magazines were on a shelf. Each book costs $5 while each magazine costs $3.
The total cost of the books was $6 more than the magazines.
How many books were on the shelf.
~~~~~~~~~~~~~~~~

x books, 30-x magazines.

Money equation is

    5x -3*(30-x) = 6  dollars.


Simplify and find x

    5x - 90 + 3x = 6

    8x = 90+6 = 96

     x        = 96/8 = 12.


ANSWER.  12 books.

Solved.